### v2

# TC-Light: Temporally Coherent Generative Rendering for Realistic World Transfer

Abstract
Illumination and texture editing are critical dimensions for world-to-world transfer, which is valuable for applications including sim2real and real2real visual data scaling up for embodied AI. Existing techniques generatively re-render the input video to realize the transfer, such as video relighting models and conditioned world generation models. Nevertheless, these models are predominantly limited to the domain of training data (e.g., portrait) or fall into the bottleneck of temporal consistency and computation efficiency, especially when the input video involves complex dynamics and long durations. In this paper, we propose TC-Light, a novel generative renderer to overcome these problems. Starting from the video preliminarily relighted by an inflated video relighting model, it optimizes appearance embedding in the first stage to align global illumination. Then it optimizes the proposed canonical video representation, i.e., Unique Video Tensor (UVT), to align fine-grained texture and lighting in the second stage. To comprehensively evaluate performance, we also establish a long and highly dynamic video benchmark. Extensive experiments show that our method enables physically plausible re-rendering results with superior temporal coherence and low computation cost. The code and video demos are available at https://dekuliutesla.github.io/tclight/.
1 Introduction
Lighting and its interaction with both real and synthetic environments fundamentally shapes how humans—and embodied agents—perceive the world. The ability to re-render the illumination and texture (or so-called relighting) of camera-captured image sequences, especially in complex, highly dynamic scenes, is critically valuable for various world-to-world transfer use cases. For instance, by re-rendering the CG-simulated or realistic video data used to train embodied agents, it can bridge the sim-to-real gap and enable real-to-real transfer, thus unlocking access to massive high-quality data that is essential for stepping towards embodied intelligence.
Despite its importance, the video illumination re-rendering remains a highly challenging problem, particularly when camera motion is highly dynamic and foreground objects frequently enter and exit scenes, which are common cases for embodied agents, as shown in Fig. 1. Most existing generative relighting techniques zhang2025scaling ; li2022physically ; jin2024neural_gaffer ; ren2024relightful ; kim2024switchlight ; wang2023sunstage are tailored for static images. As shown in Sec. 4.2, naively inflating them to a video model with existing zero-shot strategies struggles to balance the consistency and quality. Moreover, the considerable training cost and scarcity of video lighting datasets hinder fine-tuning a pretrained model for this task. Besides, though generative video relighting and world generation models are emerging, they are either restricted on domain of training data zhang2021neural ; choi2024personalized ; cai2024real ; alhaija2025cosmos or burdened by considerable computation overhead zhou2025light ; fang2025relightvid on long video, as validated in Sec. 4.2.
Figure 1: Relighting results on long videos under various dynamic scenes, averaging 256 frames per clip. Though the video involves frequent changes of foreground objects (row (a)), highly dynamic camera motions (row (b)), the TC-Light realizes consistent and physically plausible relighting results. Row (c) also shows its potential to mitigate the sim2real gap for synthetic renderings.
To address the limitations outlined above, we propose TC-Light. We utilize the SOTA image relighting model IC-Light zhang2025scaling as the baseline, and inflate it to a video model in a zero-shot manner with incorporation of our proposed decayed multi-axis denoising. It provides a preliminary video relighting result. The core innovation of TC-Light lies in a two-stage post-optimization framework that substantially improves temporal consistency. The first stage introduces per-frame appearance embedding to compensate for exposure discrepancy. It is optimized with photometric loss against the preliminarily relighted video and a flow-based loss between adjacent frames. This enforces global illumination consistency and facilitates consequent optimization. The second stage compresses the output to a canonical representation, i.e., Unique Video Tensor (UVT), according to priors including optical flow and depth of the source video. UVT is then optimized by minimizing the warping error across decompressed frames while aligning the content with the first stage result. As shown in Tab. 2, our optimization procedure is extremely efficient and introduces minimal VRAM overhead.
To comprehensively assess the effectiveness of our model, we introduce a challenging benchmark tailored for complex and highly dynamic scenes. It comprises 58 videos of averagely 256 frames per clip, spanning both indoor and outdoor environments, realistic and synthetic settings, and a wide range of lighting and weather conditions. Extensive experiments demonstrate that our method achieves high-quality, temporally consistent video relighting while maintaining low computational overhead, highlighting its strong potential for downstream applications such as embodied AI. Our main contributions are as follows:
  • $\bullet$
    A novel optimization-based video relighting paradigm for long videos with high and complicated dynamics, significantly improving the temporal consistency of the relighting result.
  • $\bullet$
    We establish a new long-video relighting benchmark characterized by high motion dynamics and broad scene diversity, covering various environments and data domains.
  • $\bullet$
    Extensive experiments validate that our method achieves SOTA performance in producing temporally consistent, naturally relighted videos with minimal computational cost.
2 Related Work
2.1 Learning-based Illumination Editing
Over the past few years, deep neural networks have become one of the main forces behind research in the field of illumination control. Pioneering works sun2019single ; nestmeyer2020learning ; pandey2021total ; das2021dsrn train convolutional encoder–decoder networks on light-stage data. The learned prior knowledge enables models to relight a portrait according to the specified light conditions. More recently, large diffusion-based generators have gained popularity for illumination editing. LightIt kocsis2024lightit explicitly conditions the diffusion process on estimated shading and normal maps, giving fine-grained lighting control ability, while SwitchLight kim2024switchlight incorporates a physics-guided architecture to simulate light-surface interactions better. zhang2024lumisculptconsistencylightingcontrol ; bharadwaj2024genlitreformulatingsingleimagerelighting leverage video foundation models to generate realistic lighting variations over a static image. IC-Light zhang2025scaling , the current state of the art, learns illumination mixture and decomposition from a large quantity of data. Building on these advances in image relighting, video relighting has started to gain traction. zhang2021neural ; choi2024personalized learns to disentangle light and intrinsic appearance on portrait videos. cai2024real represents talking faces as relightable NeRFs guided by predicted albedo and shading features. Extending IC-Light, Light-A-Video zhou2025light introduces zero-shot cross-frame attention modification, while RelightVid fang2025relightvid trains a temporally inflated IC-Light with a carefully designed video relighting dataset. However, these methods are either restricted to portrait scenarios or struggle with computational efficiency on long videos. In contrast, our model delivers high-quality relighting with strong temporal consistency and low computation cost, even in complex and highly dynamic scenes.
2.2 Diffusion-based Video Editing
The diffusion model ho2020denoising has become the go-to model for visual domain transfer and content editing. Based on training paradigms, recent advancements can be grouped into three categories: (i) training-based models extend pretrained image diffusion models with temporal layers and are trained on large-scale video datasets, such as liew2023magicedit ; chen2023controlavideo ; molad2023dreamixvideodiffusionmodels ; ma2024maskint ; qin2024instructvid2vid ; wang2025videodirector . CCEdit feng2024ccedit and FlowVid liang2024flowvid further integrate depth and flow cues for improved consistency and control. (ii) training-free models mainly rely on cross-frame attention to enforce temporal coherence. TokenFlow tokenflow2023 and FLATTEN cong2023flatten guide attention using estimated optical flow. RAVE kara2024rave enhances latent interactions by denoising over a reorganized latent grid, while Slicedit cohen2024slicedit uses spatiotemporal slices to inject motion priors. VidToMe li2024vidtome , on the other hand, exploits temporal redundancy through token merging and unmerging. (iii) one-shot-tuned models typically learn a canonical video representation in a few iterations and propagate its edits across frames. StableVideo chai2023stablevideo learns to represent video as a foreground and background atlas. CoDeF ouyang2024codef learns a hash table and decoding MLP to map frames to a single canonical image. Video-3DGS shin2024enhancing adapts deformable 3DGS kerbl3Dgaussians to model input video. Our method combines (ii) and (iii) and proposes an explicit, compact, and efficient canonical representation, i.e., Unique Video Tensor. It enables optimization to be finished within several minutes, which is much faster than 10-30 minutes cost shin2024enhancing of CoDF and Video-3DGS. Our method also inherits the diffusion model design from training-free algorithms to reduce overall memory and time cost, enabling the processing of long videos.
3 Method
3.1 Preliminaries
Task Setting. As shown in Fig. 2, we take RGB video as input. The axes of the video space-time volume are denoted by $(x,y,t)$, where $xy$ planes correspond to video frames and $yt$ planes are defined as spatiotemporal slices cohen2024slicedit . Since the camera motion is highly dynamic, the target illumination can no longer be simply appointed by a static image or an HDR environment map. Due to superior flexibility and operability, we use textual prompts as the control signal and relight the entire frame.
Latent Diffusion Models (LDMs). Denoising Diffusion Probabilistic Models (DDPMs) ho2020denoising are a class of generative models that aim to recover target data distribution through an iterative denoising process. Due to the high computational cost of operating directly in pixel space, LDMs ramesh2022hierarchical ; rombach2022high ; saharia2022photorealistic perform diffusion in a lower-dimensional latent space. Given a clean image $x_{0}$, and a pretrained autoencoder $\{\mathcal{E}(\cdot),\mathcal{D}(\cdot)\}$, LDMs first encode the image into latent space $z_{0}=\mathcal{E}(x_{0})$. The forward diffusion process then gradually corrupts $z_{0}$ with Gaussian noise $\epsilon$ over time steps $\tau=1,...,T$
$z_{\tau}=\sqrt{\alpha_{\tau}}z_{0}+\sqrt{1-\alpha_{\tau}}\epsilon,$ (1)
where $\{\alpha_{\tau}\}$ is a monotonically decreasing noise schedule. The reverse process begins from pure noise $z_{T}\sim\mathcal{N}(0,\textbf{{I}})$. With guidance from control signal (image, text, depth, etc) $c$, the trained UNet ronneberger2015u $\epsilon_{\theta}$ estimates the noise direction and progressively removes the noise from $z_{T}$. After the final denoising step, the estimated clean latent $\hat{z_{0}}$ is decoded by $\mathcal{D}(\cdot)$ to obtain the generated image $\hat{x_{0}}=\mathcal{D}(\hat{z_{0}})$, which approximates the training distribution.
Figure 2: TC-Light overview. Given the source video and text prompt $p$, the model tokenizes input latents in $xy$ plane and $yt$ plane seperately. The predicted noises are combined together for denoising (cf. Sec. 3.2). Its output then undergoes two-stage optimization to enhance temporal consistency of illumination and texture, which are respectively detailed in Sec. 3.3.1 and Sec. 3.3.2.
3.2 Lifting Image Diffusion Model to Video Space
Considering outstand ability in physical plausibility and intrinsic property preservation, we adapt IC-Light zhang2025scaling into a zero-shot video diffusion model. Concretely, we (i) enhance its diffusion blocks to capture spatiotemporal dependencies and (ii) introduce consistency prior from original frames. For (i), we apply the token merging and unmerging technique of VidToMe li2024vidtome to self-attention blocks. It divides the video frames into chunks and applies intra-chunk local token merging and inter-chunk global token merging, enabling short- and long-term consistency. The derived model $\epsilon_{\theta}$ serves as the basis of (ii). Since it reduces the token count fed to the self-attention module, the computation cost is significantly decreased. For full details, please refer to the original VidToMe paper li2024vidtome .
For (ii), we propose decayed multi-axis denoising. Similar to Slicedit cohen2024slicedit , the denoiser has two components with shared weights $\epsilon^{xy}_{\theta}(\cdot,p)$ that tokenizes each frame and merges tokens from local temporal slots, while $\epsilon^{yt}_{\theta}(\cdot,``\,")$ tokenizes the $yt$ planes (cf. Sec. 3.1) and merges tokens from local image width slot. Note that $\epsilon^{xy}_{\theta}$ conditions on target prompt $p$, while $\epsilon^{yt}_{\theta}$ takes empty prompt ``"``"``\,"` ` " as input (making the denoiser unconditional). The noises separately predicted by two parts according to the same input latents are combined together cohen2024slicedit
$\epsilon^{V}_{\theta}(\cdot,p)=\sqrt{\gamma}\epsilon^{xy}_{\theta}(\cdot,p)+\sqrt{1-\gamma}\epsilon^{yt}_{\theta}(\cdot,``\,"),$ (2)
where hyperparameter $\gamma\in[0,1]$ balances effect from $\epsilon^{yt}_{\theta}$. However, the unconditional $\epsilon^{yt}_{\theta}$ would overly biases texture and lighting toward the source video, and therefore lead to unnatural relighting results, as validated in Fig. 3 and Fig. 4. To alleviate this problem, we replace $\gamma$ with a timestep-dependent $\gamma_{\tau}$ that exponentially decays during denoising. To further align predicted noise from $\epsilon^{yt}_{\theta}$ to that of $\epsilon^{xy}_{\theta}$, we use Adaptive Instance Normalization (AIN) huang2017arbitrary to adjust feature statistics
$\epsilon_{\theta}^{V}(\cdot,p)=\sqrt{\gamma_{\tau}}\epsilon_{\theta}^{xy}(\cdot,p)+\sqrt{1-\gamma_{\tau}}\widehat{\epsilon_{\theta}^{yt}}(\cdot,``\,"),$ (3)
$\widehat{\epsilon_{\theta}^{yt}}(\cdot,``\,")=\sigma_{\epsilon_{\theta}^{xy}}\left(\frac{\epsilon_{\theta}^{yt}(\cdot,``\,")-\mu_{\epsilon_{\theta}^{yt}}}{\sigma_{\epsilon_{\theta}^{yt}}}\right)+\mu_{\epsilon_{\theta}^{xy}},$ (4)
where $\mu_{*}$ and $\sigma_{*}$ are the channel-wise mean and standard deviation of each frame. This design preserves motion guidance from the source video while reducing unwanted texture and lighting bias, as validated by ablation studies in Sec. 4.3. The output denoised video is denoted as $\{I_{t}\}$.
3.3 Post Optimization for Temporal Consistency
Although the video diffusion extension in Sec. 3.2 has introduced spatial-temporal awareness and motion prior from source video, noticeable illumination and texture flicker persist. To efficiently remove these artifacts, we introduce a two-stage post-optimization framework, as illustrated in parts (b) and (c) of Fig. 2.
3.3.1 Stage I: Exposure Alignment
As shown in part (b) of Fig. 2, the first stage introduces a per-frame appearance embedding $E_{t}$ to compensate for exposure misalignment between adjacent frames. Inspired by kerbl2024hierarchical , we model $E_{t}$ as a $3\times 4$ affine transformation matrix, initialized to the identity and optimized via Adam 2015-kingma . Its supervision combines a photometric term with a flow-warp alignment term using hyperparameter $\lambda_{e}$
$\mathcal{L}_{exposure}$ $=\left(1-\lambda_{e}\right)\mathcal{L}_{photo}\left(\tilde{I}_{t},I_{t}\right)+\lambda_{e}\mathcal{L}_{1}\left(\tilde{I}_{t}\odot M_{t},\mathrm{Warp}_{t+1\rightarrow t}\left(\tilde{I}_{t+1}\right)\odot M_{t}\right),$ (5)
where the homogeneously transformed pixel color $\tilde{I}_{t}\left(x,y\right)=E_{t}\left[I_{t}\left(x,y\right)|1\right]^{T}$. The photometric loss $\mathcal{L}_{photo}$ is the weighted sum of L1 loss and D-SSIM loss kerbl2024hierarchical , ensuring the transformed frame retains its original content and structure. The second term warps the next frame back to the current timestamp $t$, according to forward and backward flows $F_{fwd,t}$ and $F_{bwd,t}$ estimated through MemFlow dong2024memflow or provided by the dataset. Then it applies an L1 penalty $\mathcal{L}_{1}$ to align their exposures. To mask out regions with unreliable flow or occlusion, we apply a soft mask $M_{t}$
$M_{t}$ $=\mathrm{sigmoid}\left(\beta\left(\xi_{flow}-E_{flow}\right)\right)\odot\mathrm{sigmoid}\left(\beta\left(\xi_{rgb}-E_{rgb}\right)\right),$ (6)
$E_{flow}$ $=\mathrm{Norm}\left(F_{bwd,t}+\mathrm{Warp}_{t-1\rightarrow t}\left(F_{fwd,t-1}\right)\right),\quad E_{rgb}=|I_{t}-\mathrm{Warp}_{t+1\rightarrow t}\left(I_{t+1}\right)|.$ (7)
Here, $\beta$ is a constant scaling factor, $\xi_{flow}$ and $\xi_{rgb}$ are thresholds set from the statistics of error map $E_{flow}$ and $E_{rgb}$. This soft mask is also applied in the second stage of optimization. As shown in Tab. 3, soft masking outperforms the hard one in both temporal consistency and prompt alignment.
3.3.2 Stage II: Optimization over Unique Video Tensor
In the second stage, we refine illumination and texture details. Compared with vanilla video, its canonical representation can incorporate spatial-temporal priors and facilitate consistency ouyang2024codef ; shin2024enhancing . But popular NeRF or 3DGS are too complex and costly for learning (cf. Sec. 2.2). Instead, we compress the video to a one-dimensional RGB vector of shape $(N,3)$, as shown in part (c) of Fig. 2. Specifically, we define a $d$-dimensional index $\kappa(x,y,t)$ for each pixel based on priors extracted from the source video. An example index could be $[22,127,0,255]$, where the first element is flow ID (grouping pixels with shared optical flow), and the rest are 8-bit quantized RGB values. It is also allowed to extend this 4-element index to more elements with voxel coordinate (from depth projection) or any other cues that indicate spatial–temporal similarity and locality. All pixels with identical $\kappa$ are gathered via averaging to form one element of the one-dimensional vector, where $N$ is the number of unique $\kappa$. Take the source video $\{I_{t}^{in}\}$ as an example, the gathering and scattering operations are formulated as
$\mathbf{U}\left(\kappa_{n}\right)=\mathrm{Avg}\left(\left\{I_{t}^{in}\left(x,y\right)|\kappa(x,y,t)=\kappa_{n}\right\}\right),\quad\mathbb{I}_{t}^{in}\left(x,y\right)=\mathbf{U}\left(\kappa(x,y,t)\right),$ (8)
where $\mathbf{U}$ is referred to as the Unique Video Tensor (UVT). With an appropriate definition of $\kappa$, the scattered $\mathbb{I}_{t}^{in}\left(x,y\right)$ reconstructs the original $I_{t}^{in}\left(x,y\right)$ with minimal information loss, as validated in Tab. 4. For relighting, the ideal edited video frames must preserve consistent motion and intrinsic image details with the source; thus, they share the same index tensor $\kappa$ for UVT representation. Accordingly, we compress the first-stage output $\tilde{I}_{t}\left(x,y\right)$ into $\mathbf{\tilde{U}}$ via Eq. 8, which then serves as the primary optimization target. This formulation not only facilitates optimization but also naturally embeds spatial-temporal similarity priors (cf. Sec. 4.3). With CUDA parallelism, the gathering and scattering process can be performed instantly. The optimization of $\tilde{\mathbf{U}}$ is supervised by
$\mathcal{L}_{unique}=\lambda_{tv}\mathcal{L}_{tv}\left(\tilde{\mathbb{I}}_{t}\right)+\left(1-\lambda_{u}\right)\mathcal{L}_{SSIM}\left(\tilde{\mathbb{I}}_{t},\tilde{I}_{t}\right)+$ (9)
$+\lambda_{u}\mathcal{L}_{1}\left(\tilde{\mathbb{I}}_{t}\odot M_{t},\mathrm{Warp}_{t+1\rightarrow t}\left(\tilde{\mathbb{I}}_{t+1}\right)\odot M_{t}\right),$
where $\tilde{\mathbb{I}}_{t}\left(x,y\right)=\tilde{\mathbf{U}}\left(\kappa(x,y,t)\right)$, and $\lambda_{tv}$ and $\lambda_{u}\in[0,1]$ balance the loss terms. The total variation loss $\mathcal{L}_{tv}$ suppresses noise. Notably, Eq. 9 applies SSIM loss instead of photometric loss. This leaves space to fine-grained appearance and illumination adjustment without altering image structure. Finally, the optimized $\widehat{\mathbf{U}}$ is used to reconstruct $\widehat{\mathbb{I}}_{t}(x,y)$ according to Eq. 8 as the final output.
4 Experiments
4.1 Experiment Setting
Table 1: Datasets MIFDB16 ; dosovitskiy2017carla ; sun2020scalability ; Dauner2024NEURIPS ; contributors2025agibotworld ; khazatsky2024droid ; InteriorNet18 ; karnan2022scand contained in established benchmark. $N_{seq.}$ and $\bar{N}_{frames}$ denote number of sequence and average frames. C, F, D, S respectively denote RGB image, Optical Flow, Depth, Instance Segmentation. Notably, AgiBot here denotes AgiBot Digital World. Due to lacking of extrinsics, its depth is indeed not applicable. Only DRONE is self-collected data.
Datasets SceneFlow CARLA Waymo NavSim AgiBot DROID InteriorNet SCAND DRONE
Agent Vehicle Vehicle Vehicle Vehicle Robot Robot Robot Robot Drone
Synthetic
Modality C,F,D,S C,D,S C C C C C,D,S C C
$N_{seq.}$ 4 8 5 5 8 12 5 6 5
$\bar{N}_{frames}$ 300 208 198 250 305 243 300 289 213
Width 960 960 960 960 640 960 640 960 1280
Height 512 536 640 536 480 536 480 536 720
Implementation Details. Following IC-Light zhang2025scaling , we apply $T=25$ sampling steps and a classifier-free guidance scale of 2.0. When inflated to video model with VidToMe li2024vidtome , the local and global token merging ratios are 0.6 and 0.5, respectively, to accommodate high video dynamics. In our decayed multiaxis denoising strategy, the initial $\gamma_{\tau}$ is set to 0.2 and decays exponentially to 0.002 until the final sampling step. For the post-optimization stages, we use Adam kingma2017adammethodstochasticoptimization as optimizer and run 35 epochs in the first stage and 70 in the second with a batch size of 16, ensuring fast yet sufficient convergence. $\kappa(x,y,t)$ mainly contains quantized RGB and estimated masked flow, and optionally depth if provided. Emperically, the weighting coefficients $\lambda_{tv}$ is set to 0.01, $\lambda_{e}$ and $\lambda_{u}$ are set to 0.8. Following kerbl3Dgaussians , the learning rate in the first stage decays from 0.01 to 0.001, while the second stage uses a fixed learning rate of 0.05. Additional details are included in the Appendix.
Dataset. To comprehensively evaluate the generation capability, we collect video clips with high motion dynamic and broad scene diversity. This benchmark, as detailed in Tab. 1, covers scenarios like autonomous driving, robot manipulation, and navigation, as well as drone flight. It includes data collected from synthetic and realistic environments under various weather conditions. Each clip is a long video with high dynamics and on average 256 frames, making it extremely challenging. To obtain edit prompts, we use some prompts from zhang2025scaling and generate others using COSMOS agarwal2025cosmos .
Figure 3: Qualitative comparison of results. The proposed TC-Light avoids unnatural relighting like Slicedit cohen2024slicedit and COSMOS-Transfer1 alhaija2025cosmos in (a) and blurring like alhaija2025cosmos in (b), or inconsistent illumination like per-frame IC-Light zhang2025scaling and VidToMe li2024vidtome as highlighted by the red squares.
Metrics. Following prior works qi2023fatezero ; li2024vidtome ; kara2024rave ; wang2025videodirector , we assess the relighting performance along following four dimensions: (i) Temporal consistency is quantified via motion smoothness (Motion-S) huang2024vbench and structural warping error (Warp-SSIM). Motion-S evaluates the continuity and physical plausibility of motion in the edited sequence, whereas Warp-SSIM computes the SSIM between a frame and its warped neighbors using flow from RAFT teed2020raft . (ii) Textual alignment is measured by average CLIP embedding similarity between the text prompt and all edited frames (CLIP-T). (iii) User preference is evaluated by a study on 19 randomly selected videos and 65 valid submissions collected. Participants choose their preferred relighting results among our method and established baselines, from which we derive the Bradley–Terry preference rate (User-PF) bradley1952rank . Additional details are included in the Appendix. (iv) Computation efficiency is reported in terms of runtime speed (FPS) and peak GPU memory consumption (VRAM) during editing. All experiments are conducted on a 40GB A100 GPU. Additionally, to appraise the reconstruction quality of UVT, we report average PSNR, SSIM, and LPIPS between the original and reconstructed frames.
Baselines. We benchmark our approach against several recent state-of-the-art techniques, whose code is publicly available at the time of writing. These include per-frame IC-Light (denoted as IC-Light*) and its video extensions, Light-A-Video zhou2025light and RelightVid fang2025relightvid . We also implement two IC-Light variants by incorporating leading zero-shot video editing methods: VidToMe li2024vidtome and Slicedit cohen2024slicedit . For fairness, we disable the image downsampling to $512\times 512$ resolution before the diffusion step in Slicedit. In addition, we compare two advanced training-based methods—VideoDirector wang2025videodirector and COSMOS-Transfer1 alhaija2025cosmos . For the latter, due to out-of-memory (OOM) issues when applying full multimodal control on long videos, we employ only its edge branch, which offers a favorable balance between preserving image details and adhering to relighting prompts.
4.2 Comparison with SOTA
Table 2: Comparison with existing methods. "OOM" here means the method is unable to finish the task due to an out-of-memory error. For a fair comparison, the base models of VidToMe and Slicedit are replaced with IC-Light here. Ours-light applies post-optimization to VidToMe, while Ours-full further introduces decayed multi-axis denoising. Experiments are conducted on 40G A100. The best and the second best of each metric are separately highlighted in red and blue.
Method Motion-S↑ WarpSSIM↑ CLIP-T↑ User-PF↑ FPS↑ Time(s)↓ VRAM(G)↓
IC-Light* zhang2025scaling 94.52% 71.22 0.2743 10.97% 0.123 2075 16.49
VidToMe li2024vidtome 95.38% 73.69 0.2731 6.97% 0.409 626 11.65
Slicedit cohen2024slicedit 96.48% 85.37 0.2653 18.39% 0.122 2101 17.87
VideoDirector wang2025videodirector OOM OOM OOM OOM OOM OOM OOM
Light-A-Video zhou2025light OOM OOM OOM OOM OOM OOM OOM
RelightVid fang2025relightvid OOM OOM OOM OOM OOM OOM OOM
Cosmos-T1 alhaija2025cosmos 96.83% 83.47 0.2529 16.06% 0.101 2543 34.87
Ours-light 97.39% 88.53 0.2700 23.66% 0.359 771 14.36
Ours-full 97.80% 91.75 0.2679 23.96% 0.204 1255 14.37
Quantitative and qualitative comparisons with state-of-the-art methods are reported in Tab. 2 and Fig. 3, respectively. The result indicates that per-frame relighting (IC-Light*) follows prompts well and produces physically plausible illumination, but the adapted illumination suffers from severe flicker, as shown in columns (a) and (b) of Fig. 3. IC-Light would even randomly hallucinate non-existent objects in textureless regions (cf. column (c) of Fig. 3), further degrading consistency. Extending IC-Light* with VidToMe li2024vidtome yields modest gains in temporal coherence but dramatically lowers computation cost for long videos, so we adopt it as our primary baseline. Slicedit cohen2024slicedit significantly suppresses flicker and hallucinations, yet its computation overhead exceeds that of IC-Light*. Besides, its output remains overly biased by the original appearance of the source video. As a result, it produces unnatural relighting in many cases, as shown in column (a) of Fig. 3.
We also evaluated the T2V-model-based video editing approach wang2025videodirector and concurrent video relighting techniques zhou2025light ; fang2025relightvid . Unfortunately, they all failed on long clips due to OOM errors caused by high computation resource demands. For the same reason, Cosmos-Transfer1 alhaija2025cosmos can only operate in single-modality mode under GPU constraints, yet still requires over 30 GB GPU memory and more than 30 minutes per clip. Moreover, on video with high dynamics, it suffers from more severe blur and loss of details, as shown in columns (a) and (b) of Fig. 3. These failures are likely because Cosmos-Transfer1 is limited to the data domain of its training data, which contains less varied, moderately dynamic videos.
In contrast, our TC-Light first enables physically plausible relighting on long videos with high dynamics, while outperforming all baselines in temporal consistency and preference rate by a large margin, as shown in Tab. 2. The light version adds only 2.4 minutes and 2.7 GB of VRAM overhead compared to the VidToMe baseline, while faithfully preserving object identity, albedo, and adherence to text prompts, as shown in Fig. 3. Incorporating our decayed multi-axis denoising further enhances temporal coherence, with a modest trade-off in efficiency and quality. Limited by page, we provide additional visualization and performance of different scenarios types in the Appendix.
4.3 Ablation
Table 3: Ablation over module component. The experiments here are conducted on CARLA dosovitskiy2017carla and the Interiornet InteriorNet18 subset, which both provide depth and instance mask as priors. There are 13 sequences in total and 254 frames on average, covering scenes of indoor and outdoor scenarios. The gray row denotes modification that is aborted and not included in the following experiments.
Method Motion-S↑ WarpSSIM↑ CLIP-T↑ FPS↑ Time(s)↓ VRAM(G)↓
Baseline 94.51% 77.60 0.2871 0.693 364 10.63
+1st Stage 95.71% 81.29 0.2868 0.651 388 11.33
+2nd Stage(video) 96.40% 90.58 0.2876 0.552 460 13.53
+2nd Stage(UVT) 96.44% 91.04 0.2866 0.563 449 11.81
+soft mask 96.44% 91.05 0.2868 0.559 452 11.81
from scratch(UVT) 96.30% 90.65 0.2866 0.552 458 12.40
+depth 96.56% 91.12 0.2863 0.569 444 11.57
+instance 96.50% 91.01 0.2851 0.545 462 11.67
+multi-axis 98.41% 95.52 0.2813 0.310 805 11.57
+AIN 98.38% 95.44 0.2832 0.310 805 11.57
+weight decay 97.75% 93.74 0.2865 0.310 805 11.57
Table 4: Ablation over Unique Video Tensor (UVT). Here, %$\mathrm{Cmpr}$ is the compression rate after applying UVT on the source video. The subscripts “f” and “f+d” indicate that, besides color cues, the UVT representation incorporates optical flow cues and both flow and depth cues, respectively.
Scene %$\mathrm{Cmpr}_{f}$ $\mathrm{SSIM}_{f}$ $\mathrm{PSNR}_{f}$ $\mathrm{LPIPS}_{f}$$\mathrm{Cmpr}_{f+d}$ $\mathrm{SSIM}_{f+d}$ $\mathrm{PSNR}_{f+d}$ $\mathrm{LPIPS}_{f+d}$
CARLA %39.2 0.9940 50.71 0.025 %29.2 0.9925 48.98 0.028
InteriorNet %49.0 0.9908 46.17 0.021 %12.8 0.9755 40.86 0.047
Figure 4: Ablation on main module components. The experiment is conducted on one sequence of the InteriorNet InteriorNet18 subset, where the text prompt is "This video showcases a modern interior space, which is dimly lit". The baseline here denotes VidToMe li2024vidtome in Tab. 2.
This section analyzes the contribution of each component in our model. The first stage optimization, as shown in Tab. 3 and Fig. 4, markedly boosts consistency by aligning cross-frame exposure. The 6-7th rows of Tab. 3 also illustrate that, initializing UVT optimization from the first-stage results converges more efficiently than directly optimizing UVT for the same overall epochs. The second stage optimization, as shown in Tab. 3, further reinforces temporal coherence. Tab. 4 confirms that UVT can compress the source video with near-zero loss, which underpins our design in Sec. 3.3.2. Using UVT as the second-stage target not only boosts consistency but also cuts computational overhead. Additionally, replacing a hard mask with a soft mask consistently improves both Warp-SSIM and CLIP-T metrics, demonstrating its importance. Incorporating the depth cues alongside UVT yields a more compact representation (also in Tab. 4), which aids illumination alignment and release computation burden. In contrast, instance segmentation masks provide no clear benefit and are thus omitted from the final implementation.
For the diffusion module, multi-axis denoising notably enhances temporal consistency. However, it tends to inherit appearance distribution from the source video, causing drift from the target prompt and sometimes unnatural lighting, as shown in Fig. 4. The introduced AIN and weight decay mitigate these issues, achieving a promising balance between consistency and faithful prompt alignment.
4.4 Limitation and Discussion
Despite achieving impressive results, our method is still limited by its base models. For instance, the current version of IC-Light zhang2025scaling still struggles to relight hard shadows or make large modifications to low-light images. Similarly, since IC-Light is pretrained on 512 resolution and fine-tuned on 1024 resolution, our model struggles to preserve image details if the resolution is lower than 512. Besides, since the optimization process relies on the optical flow estimation model, artifacts sometimes occurs in textureless areas where flow becomes unreliable. Furthermore, the temporal consistency loss has the tendency to smooth the texture of flickering areas, and therefore might sacrifice some details. Though the proposed decayed multiaxis denoising alleviates the problem, developing a temporally more consistent and computationally more efficient denoising strategy is desired in future work.
5 Conclusion
In summary, we present TC-Light, a one-shot-tuned framework that delivers temporally consistent and physically plausible relighting on long, highly dynamic videos. The optimization-based illumination alignment provides a new paradigm for video relighting. Central to our approach is the Unique Video Tensor—an explicit, canonical, and differentiable video representation that enables highly efficient optimization. Over the established long video relighting benchmark, TC-Light achieves state-of-the-art performance in both consistency and efficiency, making it particularly well suited for sim2real and real2real data scaling in embodied AI.
Appendix A Additional Experimental Results
Figure 5: Qualitative results on additional long highly dynamic videos.
Figure 6: Additional qualitative comparison of results. The proposed TC-Light avoids unnatural relighting like Slicedit cohen2024slicedit and COSMOS-Transfer1 alhaija2025cosmos in (a) and (b), or temporal inconsistency like per-frame IC-Light zhang2025scaling and VidToMe li2024vidtome as highlighted by the red squares.
Table 5: Comparison on synthetic MIFDB16 ; dosovitskiy2017carla ; contributors2025agibotworld ; InteriorNet18 and realistic scenarios sun2020scalability ; Dauner2024NEURIPS ; khazatsky2024droid ; karnan2022scand . The average resolutions are respectively $794\times 503$ and $960\times 555$, while the frame numbers are 272 and 246. "OOM" here means the method is unable to finish the task due to an out-of-memory error. Ours-light applies post-optimization to VidToMe, while Ours-full further introduces decayed multi-axis denoising. The best and the second best of each metric are separately highlighted in red and blue.
Synthetic Realistic
Method Motion-S↑ WarpSSIM↑ CLIP-T↑ Motion-S↑ WarpSSIM↑ CLIP-T↑
IC-Light* zhang2025scaling 93.43% 66.02 0.2779 95.14% 77.13 0.2837
VidToMe li2024vidtome 94.61% 69.45 0.2776 95.82% 79.33 0.2815
Slicedit cohen2024slicedit 96.28% 84.90 0.2717 96.38% 88.89 0.2715
VideoDirector wang2025videodirector OOM OOM OOM OOM OOM OOM
Light-A-Video zhou2025light OOM OOM OOM OOM OOM OOM
RelightVid fang2025relightvid OOM OOM OOM OOM OOM OOM
Cosmos-T1 alhaija2025cosmos 96.31% 80.87 0.2537 96.78% 83.57 0.2659
Ours-light 97.02% 88.63 0.2707 97.46% 89.42 0.2816
Ours-full 97.36% 91.07 0.2695 97.90% 92.67 0.2792
Fig. 5 presents additional visualizations of our relighting results across a diverse range of scenarios. Whether under nighttime or daytime conditions, in outdoor or indoor environments, or from aerial or ground-level viewpoints, the proposed TC-Light method consistently produces temporally coherent and physically plausible illumination edits, demonstrating strong generalization capabilities. Fig. 6 provides qualitative comparisons against state-of-the-art methods across additional scenarios. As shown, our model effectively adheres to textual instructions while generating relighting results that are both natural and temporally consistent.
We also provide corresponding quantitative evaluations on synthetic and real-world scenarios. As reported in Tab. 5, performance on real-world scenes consistently exceeds that on synthetic ones. This discrepancy likely arises from the training data of the video model Cosmos-Transfer1 alhaija2025cosmos and the foundational image model IC-Light zhang2025scaling , which are biased towards realistic scenes. Furthermore, the higher resolution and richer textures of real-world data mitigate hallucinations in textureless regions and help better preserve the intrinsic details of source frames for IC-Light. Such attributes are particularly critical for the consistency of methods with comparatively limited temporal modeling, namely, IC-Light* and VidToMe, which exhibit substantially higher Motion-S and WarpSSIM metrics on real-world videos than on synthetic ones. In contrast, our approach attains state-of-the-art temporal consistency across both scenario types while maintaining a favorable balance with prompt adherence.
Appendix B Details of Assets
Table 6: Licenses and video resolution of datasets MIFDB16 ; dosovitskiy2017carla ; sun2020scalability ; Dauner2024NEURIPS ; contributors2025agibotworld ; khazatsky2024droid ; InteriorNet18 ; karnan2022scand contained in established benchmark. Notably, AgiBot here denotes AgiBot Digital World. DRONE is our self-collected subset. Sceneflow has no license, but is only allowed for research purposes.
Datasets SceneFlow CARLA Waymo NavSim AgiBot DROID InteriorNet SCAND DRONE
Width 960 960 960 960 640 960 640 960 1280
Height 512 536 640 536 480 536 480 536 720
License N/A CC-BY Custom [^1] https://waymo.com/open/terms/ CC BY- NC-SA 4.0 CC BY- NC-SA 4.0 Apache -2.0 Custom [^2] https://interiornet.org/ CC0 1.0 N/A
In Tab. 6, we summarize the license and resolution for each subset. All source videos are resized and center-cropped to their designated resolutions. Considering the computation source limitation, we sample videos that contain over 600 frames to around 300 frames. The DRONE subset includes three clips captured using our DJI Mini4 Pro and two additional clips obtained from DroneStock [^3] https://dronestock.com/, which are released under the CC0 1.0 License. For AgiBot Digital World contributors2025agibotworld , where the robot’s head moves in coordination with its body while performing tasks, relighting is performed from the head-mounted camera view. For each scene of DROID khazatsky2024droid , we apply relighting to both the static side camera and the dynamic left wrist camera views. For Waymo sun2020scalability and NavSim Dauner2024NEURIPS , relighting is conducted using the front-facing camera view.
This paper also benefits from the code of IC-Light zhang2025scaling (Apache-2.0 License), VidToMe li2024vidtome (MIT License), Slicedit cohen2024slicedit (MIT License), VideoDirector wang2025videodirector (MIT License), Light-A-Video zhou2025light (Apache-2.0 License), RelightVid fang2025relightvid (CC BY-NC-SA 4.0 License), and Cosmos-Transfer1 alhaija2025cosmos (Apache-2.0 License).
Appendix C Additional Implementation Details
For competing methods, we adopt the hyperparameters from their official implementations for VideoDirector wang2025videodirector , Light-A-Video zhou2025light , RelightVid fang2025relightvid , and Cosmos-T1 alhaija2025cosmos . We replace base models of VidToMe li2024vidtome and Slicedit cohen2024slicedit with IC-Light zhang2025scaling , and therefore we align their classifier-free guidance scale and diffusion sampling steps with those in zhang2025scaling . Additionally, we set VidToMe’s local and global token-merging ratios to 0.6 and 0.5, respectively, mirroring the setting of our approach. For Slicedit, we adjust the weighting factor $\gamma$ in Eq. (2) from the default 0.2 to 0.05 to better balance temporal coherence and instruction adherence. All other hyperparameters remain at their default values. The modified VidToMe serves as the baseline of our model design.
During implementation, each $\kappa(x,y,t)$ comprises three components: (1) a per-pixel flow ID, (2) a quantized RGB color, and, optionally, (3) a world-frame voxel coordinate. For (1), flow IDs are derived from the optical flow estimated by the state-of-the-art MemFlow method dong2024memflow and the binary mask obtained by thresholding the soft mask in Eq. (6) of the main paper (values > 0.5 are set to 1; otherwise 0). In the initial frame, pixels receive unique flow IDs from 00 to $HW-1$, where $H$ and $W$ denote image height and width. In subsequent frames, a pixel inherits the flow ID of its predecessor if connected by an unmasked flow; otherwise, it is assigned a new ID. This injects motion priors into the UVT representation. For (2), we quantize RGB values to 7 bits, ensuring that all pixels sharing the same UVT element differ by less than $2/255$ in any channel. This constraint mitigates erroneous flows that escape the mask and reinforces representation in regions exhibiting view-dependent effects. For (3), when per-frame depth maps are available, they are reprojected into a point cloud using the camera intrinsics and extrinsics to determine world-frame coordinates. This point cloud is then voxelized at a specified voxel size, and each pixel’s voxel coordinate is appended to $\kappa(x,y,t)$, yielding a more compact representation of static regions. For the CARLA dosovitskiy2017carla and InteriorNet InteriorNet18 datasets, voxel sizes are set to 0.05 m and 0.02 m, respectively. Notably, dynamic objects at different timestamps may spatially overlap in 3D, but they remain distinguishable by their flow IDs and quantized RGB colors. Consequently, each object at each timestep is represented by a distinct set of UVT elements, while the $\mathcal{L}_{1}$ temporal consistency loss preserves object identity across frames.
Appendix D User Study
Figure 7: A screenshot of the user study.
We conducted an online user study with 78 anonymous participants, evaluating 19 randomly selected video–text pairs from our datasets. The compared methods were IC-Light* zhang2025scaling , VidToMe li2024vidtome , Slicedit cohen2024slicedit , Cosmos-Transfer1 alhaija2025cosmos , Ours-light, and Ours-full. A screenshot of the questionnaire interface is shown in Fig. 7. For each question, methods were anonymized and relighted videos were presented in random order; participants selected the two most preferred results. In compliance with the NeurIPS Code of Ethics, each participant received a compensation of \$0.70. Besides, we ensured that all collected data remained confidential and was not disclosed to any institutions or individuals.
Since each video spanned 10–20 seconds, completing the questionnaire took on average 13.5 minutes. Submissions requiring less than four minutes were deemed unreliable and excluded, yielding 65 valid responses. Fig. 8 reports the frequency with which each method was chosen among the top two. Our full model achieved the highest preference rate, while the light variant ranked second. Although IC-Light* and VidToMe follow instructions well (cf. Tab. 2 of the main paper), their inferior temporal consistency make them much less preferred by users. Finally, we computed Bradley–Terry preference scores bradley1952rank as a comprehensive metric of user preference, as presented in the Tab. 2 of the main paper.
Figure 8: Results from user study with 65 valid submissions. The methods are arranged in alphabetical order. This figure reports the frequency that each method is chosen as the first- and second-most preferred video.
Appendix E Social Impact
Positive Impacts. The proposed TC-Light framework for long video relighting stands to benefit a wide range of applications in both industry and research. First, by enabling consistent and physically plausible illumination editing at low computational cost, it can substantially lower the barrier to high-quality visual content creation, empowering independent filmmakers, educators, and artists to produce compelling video narratives without access to specialized hardware. Second, the capability to scale illumination-diverse training data through sim2real and real2real transfer can accelerate progress in embodied AI—robots and autonomous agents exposed to rich, temporally coherent visual environments may learn more robust perception and planning behaviors, thereby advancing safety and reliability in human–robot interaction. Finally, by fostering more efficient video synthesis pipelines, TC-Light may encourage energy-aware design practices in large-scale media processing systems, contributing to reduced resource consumption and attendant carbon emissions.
Negative Impacts. Despite these benefits, improved video relighting carries potential risks if misused. Enhanced realism in dynamic relighting could facilitate the creation of deceptive multimedia, including deepfake videos that manipulate shadows and highlights to conceal tampering or impersonate individuals, thereby eroding trust in digital media. Moreover, large-scale deployment of relighting tools raises privacy concerns: adversarial actors might relight surveillance footage to obscure identities or fabricate altered event sequences. To mitigate these harms, we advocate for gated access to pretrained models, integration of provenance metadata to flag relit content, and collaboration with platform providers to monitor and throttle suspicious bulk relighting requests.